• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 04:33:28 on June18,2010

    Entry number 327199


    keyword=investigating R-HRS problem (low rate)

    Problem: R-HRS rate is very low in det1,2.

    Will try to diagnose with low-current T5. BTW, the last run
    27906 had ps8=2.

    Laser 65, slit 13.9
    L-HRS T5 rate = 17.2 kHz
    R-HRS T5 rate = 95 Hz

    04:10 .... VDCs turned on.

    Do a run 27909 and 6956.

    Problem: Cannot see much diff between beam "on" and "off" on L-RHS.
    Rates about 12 kHz vs 11.5 khz

    Ask for slit = 15.
    This gives 159 kHz (Left) and 5.8 kHz (Right).

    Runs 27911 and 6957.

    Comparing to earlier alignment runs, the Y looks ok (not shown)
    but the X seems off by 1 cm (see fig 1). We aren't in counting
    mode for Quartz.

    FIGURE 1